Really need help with this, thanks. Round to 2 decimal places

"Derek plans to buy a $33,451.00 car. The dealership offers zero percent financing for 56.00 months with the first payment due at signing (today). Derek would be willing to pay for the car in full today if the dealership offers him $____ cash back. He can borrow money from his bank at an interest rate of 4.63%."
Answer #1

Hi

At first we will find out monthly payment using excel PMT function

here RATE = 0%

number of period nper = 56

Present Value PV = $33,451

monthly payment = PMT(RATE,nper, pv,fv)

=PMT(0%,56,33451,1) = $597.36

Now we will find out present value based on 4.63% borrowing rate.

monthly rate RATE= 4.63%/12

    nper = 56 months

monthly payment PMT = $597.36

Present Value PV = PV(rate,nper,pmt,fv)

=PV(4.63%/12,56,597.36,0)

= $30,033.01

So cash back required = 33451- 30,033.01

= $3,417.99

Thanks